# Semiparametric conformal prediction

### v1

@@ Abstract
Many risk-sensitive applications require well-calibrated prediction sets over multiple, potentially correlated target variables, for which the prediction algorithm may report correlated non-conformity scores. In this work, we treat the scores as random vectors and aim to construct the prediction set accounting for their joint correlation structure. Drawing from the rich literature on multivariate quantiles and semiparametric statistics, we propose an algorithm to estimate the $1-\alpha$ quantile of the scores, where $\alpha$ is the user-specified miscoverage rate. In particular, we flexibly estimate the joint cumulative distribution function (CDF) of the scores using nonparametric vine copulas and improve the asymptotic efficiency of the quantile estimate using its influence function. The vine decomposition allows our method to scale well to a large number of targets. We report desired coverage and competitive efficiency on a range of real-world regression problems, including those with missing-at-random labels in the calibration set.
@@ 1 INTRODUCTION
Many problems in industrial and pharmaceutical applications can be framed as multi-target regression, where the goal is to predict a vector target taking values in $\mathbb{R}^{d}$ from some input features. In early-stage drug discovery, for example, it is necessary to jointly characterize the ADME (absorption, distribution, metabolism, and excretion) properties of a candidate molecule. The complex biophysical processes governing these properties manifest as correlations in their measurements, often pronounced in the tails [1, 2, 3].
@@
Figure 1: Top: Single-target split CP. Green dashed line in (a) marks the $1-\alpha$ quantile of calibration scores, setting the conf. interval around test predictions in (b). Bottom: Independent split CP for two targets, with $\sqrt{1-\alpha}$ quantiles marked by green dashed lines in (c). Large pred. sets for two test instances shown in (d).
@@
We often require prediction sets capturing uncertainties, rather than point estimates [4, 5]. The medicinal chemists may use the uncertainties on a given molecule to decide whether to test it in the wet lab. One way to express uncertainties is as confidence at a specified coverage level. Conformal prediction (CP) stands out as a powerful framework that can wrap around any underlying prediction algorithm to produce valid confidence sets for any data distribution [6, 7, 8]. A confidence set at the $1-\alpha$ level for a random variable $Y$, denoted ${\Gamma}_{1-\alpha}$, is valid if it contains the true value of $Y$ at least $1-\alpha$ of the time, that is, $\mathbb{P}[Y\in{\Gamma}_{1-\alpha}]\geq 1-\alpha$.
@@
Central to the CP algorithm is the concept of ranks and quantiles [9]. The confidence set at level $1-\alpha$ is constructed by extracting the $1-\alpha$ quantile of the non-conformity score quantifying how much a prediction deviates from the labels. In the context of split CP [10, 11] for single-target regression, for instance, one may define the score as the absolute error and rank the scores evaluated on the calibration set to obtain the $1-\alpha$ quantile. See Figure 1(a, b) for an illustration.
@@
CP can be applied independently for each target, not accounting for the correlation in the scores [10]. Ranking then proceeds like the single-target case at a marginal level $1-\alpha_{j}>1-\alpha$ for each target $j$ that would satisfy the desired global coverage level of $1-\alpha$. By treating the scores as independent, however, we miss out on potential gains in efficiency, or the tightness of the prediction [12]. Figure 1(c, d) illustrates independent CP for two-target regression. Setting $1-\alpha_{j}=\sqrt{1-\alpha}$ marginally results in large confidence sets. In general, the calibration set may not be large enough to accommodate extreme $1-\alpha^{\prime}_{j}$ levels, forced to be closer to 1 as $d$ increases.
@@
We instead model the targets jointly by viewing the target-wise scores as random vectors $S$ that take values in $\mathbb{R}^{d}$. The key question then is how to define the $1-\alpha$ quantile of observed $S$. While ranks and quantiles in $\mathbb{R}$ form the basis of nonparametric statistics [13], a canonical ordering does not exist in $\mathbb{R}^{d}$ for $d>1$. Several notions of multivariate quantiles have been proposed, including ones based on statistical depth [14, 15] and geometry [16, 17]. We build on the work of Messoudi et al. [18] that proposed using the copula, a statistical tool for modeling the dependence structure of multivariate random variables [19]. Copulas can be used to estimate the joint cumulative distribution function (CDF) of the scores, $F(s)=\mathbb{P}[S_{1}\leq s_{1},\dotsc,S_{d}\leq s_{d}]$, where $s\in\mathbb{R}^{d}$. The quantile is its generalized inverse, or the $p$-level curve $F^{-1}(p)={\rm inf}\{s\in\mathbb{R}^{d}:F(s)\geq p\}$. Figure 2 depicts the level curves of the copula (colored contours) fit on correlated scores (gray dots). For $0.91-\alpha=$ 0.9, the $0.9$-level line contains quantiles with smaller norms compared to the independent target approach (green dot), indicating that joint modeling can yield more efficient prediction sets.
@@
As an update to prior work, we estimate $F$ using vine copulas [20], specifically the nonparametric variant [21] for additional expressibility. Vine copulas can scale well to many targets, as they factorize high-dimensional densities into a product of bivariate densities. Applications of copulas for CP have previously been limited to the empirical copula or a small set of parametric families, which can be biased when observations are sparse or the model misspecified [18, 22]. We also demonstrate that a particular copula model allows working with missing-at-random (MAR) target labels in the calibration set.
Having inferred $F$, rather than simply using the plug-in estimate of the quantile $F^{-1}(1-\alpha)$, we tailor our analysis to this specific functional by computing the one-step estimator [23] based on its influence function [24], best understood as a generalized gradient of the functional with respect to small perturbations of $F$. This debiased estimator specifically targets the quantile estimand of interest and carries better theoretical properties like $\sqrt{n}$-consistency and asymptotic normality [25].
Contributions: (1) We propose the use of nonparametric vine copulas for flexibly estimating the joint distribution of the non-conformity scores in multivariate conformal calibration (Section 5.1). (2) We propose the semiparametric one-step estimator for debiasing the $1{-}\alpha$ score quantile used to construct the prediction sets (Section 5.2). (3) We empirically demonstrate the validity and efficiency of our method on various real-world regression problems (Section 6.2).
@@ 2 RELATED WORK
A large body of work aims at extending conformalized quantile regression [CQR; 26] to multiple targets [27, 28, 29]. There are parallel efforts to make full CP more computationally efficient for some classes of multi-target predictiors [30, 31]. Our focus is to improve the calibration step without additional assumptions on the underlying algorithm.
When there are multiple response variables, a common approach is to design a single scalar score, such as the $L_{2}$ distance between the vector-valued prediction and the label [32, 33, 34, 29]. This implies choosing a target aggregation scheme, however, and the resulting score is sensitive to the relative scales in the predictions and labels across targets. Moreover, the labels for all $d$ targets need to be observed for score evaluation, which may be impractical for large $d$.
Another approach is to conformalize the predictions for each individual target [12]. Non-conformity scores are evaluated target-wise and the score at the $1-\alpha_{j}$ quantile extracted for each target $j$ such that $\prod_{j}(1-\alpha_{j})=1-\alpha$. If all targets are to be treated equally, we set $1-\alpha_{j}=(1-\alpha)^{1/d}>1-\alpha$ for all $j$. Note that as $d$ increases, $(1-\alpha)^{1/d}$ approaches $1$, leading to inefficient prediction sets. Such extreme levels also require the calibration dataset to be sufficiently large.
If the scores are highly correlated, we can exploit the dependence structure to improve efficiency. Messoudi et al. [18] pioneered the use of copulas for jointly modeling the distribution of vector scores evaluated target-wise. Sun and Yu [22] applied this idea to multi-step time series forecasting. The empirical copulas used in previous work can be unstable with sparse observations, especially in the upper tails near the quantile level of interest. Messoudi et al. [18] also uses simple parametric copulas like Gumbel and Gaussian, which offer a smoother fit but can suffer from model misspecification. Copulas are reviewed in Section 4.
@@ 3 Problem setup
A set is marginally valid [^1] This is a weaker condition than conditional validity, $\mathbb{P}[Y^{*}\in{\Gamma}_{1-\alpha}(X^{*})|X^{*}]\geq 1-\alpha$. if it contains the response $Y^{*}$ with probability at least $1-\alpha$:
@@ 4 BACKGROUND
constructs prediction sets without any assumptions on the data distribution $P_{XY}$ [8]. The set is guaranteed to be marginally valid for finite samples as long as the data are exchangeable. CP was originally formulated in the transductive setting [6, 7, 8]. Here we focus on split (inductive) CP [10, 35], where the labeled dataset $\{(X^{(i)},Y^{(i)})\}_{i\in\mathcal{I}}$ with index set $\mathcal{I}$ is split into the proper training data $\mathcal{I}_{\rm train}$, used to fit the underlying algorithm ${\hat{f}}$, and the calibration data $\mathcal{I}_{\rm cal}$, used to conformalize it. [^2] We present our framework in terms of split CP for clarity, but it applies to full (transductive) [8] and CV [36], jackknife [37] variants as well.
@@
where ${\hat{F}}(s)=\frac{1}{n}\sum_{i\in\mathcal{I}_{\rm cal}}\mathbbm{1}[S^{(i)}\leq s]$ is the empirical CDF. Applying Equation 2 yields the confidence interval $[{\hat{f}}(X^{*})-Q_{1-\alpha},{\hat{f}}(X^{*})+Q_{1-\alpha}]$. We refer readers to [38, 39] for more details.
@@
Given a random vector $S=[S_{1},\dotsc,S_{d}]$, denote the marginal CDF for each component variable $S_{j}$, $j\in[d]$, as $F_{j}(s_{j})=\mathbb{P}[S_{j}\leq s_{j}]$. The copula $C:[0,1]\rightarrow[0,1]$ of $S$ is the joint CDF of $U=[F_{1}(S_{1}),\dotsc,F_{d}(S_{d})]$:
@@
Theorem 1 (Sklar’s theorem [40]).
@@
The vine copula, or the pair copula construction, can accommodate dependencies of arbitrary complexity when a single standard copula model is inadequate. It is a hierarchical model composed of bivariate copula blocks organized in a sequence of nested trees, also called regular (R-) vines or simply vines [41]. The main idea is to factorize a $d$-dimensional copula density into a product of ${d(d-1)}/{2}$ bivariate conditional copula densities. Fitting a vine then boils down to choosing the bivariate copula models and the vine structure. The bivariate copulas can be estimated in a nonparametric manner [42, 43, 21]. See Section 9 for more details on nonparametric vine copula estimation.
@@
refers to the inference of nonparametric estimands, or functionals of the true data distribution $P$. Also called targeted or debiased learning, this approach has emerged as a major paradigm in recent years [44]. If the target functional of interest is “smooth” (in ways we will make concrete below), it is sometimes possible to endow estimators with parametric properties like $\sqrt{n}$-consistency and asymptotic normality, even if our estimate of $P$ converges at a slower rate [45, 44, 46, 25].
@@
The second equality follows from restricting ourselves to mean-zero functions for $\psi_{P}$, as the middle expression is insensitive to constant shifts in $\psi_{P}$. When Equation 10 exists and is finite for any “regular” parametric submodel (for which the score $\tilde{P}(Z)/P(Z)-1$ has finite variance), $\psi_{P}$ is called the efficient influence function (EIF) and $\Psi$ is said to be pathwise differentiable [47]. The EIF serves as a first-order distributional derivative that characterizes how sensitive $\Psi(P)$ is to changes in $P$ [48, 46].
@@
The one-step estimator [23, 49, 50] is
@@
where $r_{2}$ is the remainder term [51]. Applying the Riesz representation theorem as in Equation 10 gives
@@
Denoting the estimation error of $\hat{P}$ as $\varepsilon$, we might expect $r_{2}\sim\mathcal{O}(\varepsilon^{2})$. This suggests $\Psi({\hat{P}})+\mathbb{E}_{P}[\psi_{\hat{P}}]$ would be better than the plug-in estimator $\Psi({\hat{P}})$. But $P$ is unknown, of course, so we take the expectation over $\mathbb{P}_{n}$. Figure 3 helps interpret one-step correction as a linear projection [52]. See Section 11 for a self-contained introduction to one-step estimation.

### v2

@@ Abstract
Many risk-sensitive applications require well-calibrated prediction sets over multiple, potentially correlated target variables, for which the prediction algorithm may report correlated errors. In this work, we aim to construct the conformal prediction set accounting for the joint correlation structure of the vector-valued non-conformity scores. Drawing from the rich literature on multivariate quantiles and semiparametric statistics, we propose an algorithm to estimate the $1-\alpha$ quantile of the scores, where $\alpha$ is the user-specified miscoverage rate. In particular, we flexibly estimate the joint cumulative distribution function (CDF) of the scores using nonparametric vine copulas and improve the asymptotic efficiency of the quantile estimate using its influence function. The vine decomposition allows our method to scale well to a large number of targets. As well as guaranteeing asymptotically exact coverage, our method yields desired coverage and competitive efficiency on a range of real-world regression problems, including those with missing-at-random labels in the calibration set.
Keywords conformal prediction  $\cdot$ nonparametric statistics  $\cdot$ copulas
@@ 1 INTRODUCTION
Many problems in industrial and pharmaceutical applications can be framed as multi-target regression, where the goal is to predict a vector target taking values in $\mathbb{R}^{d}$ from some input features. In early-stage drug discovery, for example, it is necessary to jointly characterize the ADME (absorption, distribution, metabolism, and excretion) properties of a candidate molecule. The complex biophysical processes governing these properties manifest as correlations in their measurements, often pronounced in the tails (Jain et al.,, 2017; Niepel et al.,, 2019; Du et al.,, 2023).
@@
Figure 1: Top: Single-target split CP. Green dashed line in (a) marks the $1-\alpha$ quantile of calibration scores, setting the confidence interval around test predictions in (b). Bottom: Independent split CP for two targets, with $\sqrt{1-\alpha}$ quantiles marked by green dashed lines in (c). Large prediction sets for two test instances shown in (d).
@@
We often require prediction sets capturing uncertainties, rather than point estimates (Eklund et al.,, 2015; Cortes-Ciriano and Bender,, 2020). The medicinal chemists may use the uncertainties on a given molecule to decide whether to test it in the wet lab. One way to express uncertainties is as confidence at a specified coverage level. Conformal prediction (CP) stands out as a powerful framework that can wrap around any underlying prediction algorithm to produce valid confidence sets for any data distribution (Gammerman et al.,, 1998; Saunders et al.,, 1999; Vovk et al.,, 2005). A confidence set at the $1-\alpha$ level for a random variable $Y$, denoted ${\Gamma}_{1-\alpha}$, is valid if it contains the true value of $Y$ at least $1-\alpha$ of the time, that is, $\mathbb{P}[Y\in{\Gamma}_{1-\alpha}]\geq 1-\alpha$.
@@
Central to the CP algorithm is the concept of ranks and quantiles (Kuchibhotla,, 2020). The confidence set at level $1-\alpha$ is constructed by extracting the $1-\alpha$ quantile of the non-conformity score quantifying how much a prediction deviates from the labels. In the context of split CP (Papadopoulos et al.,, 2002; Lei et al.,, 2015) for single-target regression, for instance, one may define the score as the absolute error and rank the scores evaluated on the calibration set to obtain the $1-\alpha$ quantile, as illustrated in Figure 1(a, b).
@@
CP can be applied independently for each target, not accounting for the correlation in the scores (Papadopoulos et al.,, 2002). Ranking then proceeds like the single-target case at a marginal level $1-\alpha_{j}>1-\alpha$ for each target $j$ that would satisfy the desired global coverage level of $1-\alpha$. By treating the scores as independent, however, we miss out on potential gains in efficiency, or the tightness of the prediction (Neeven and Smirnov,, 2018). Figure 1(c, d) illustrates independent CP for two-target regression. Setting $1-\alpha_{j}=\sqrt{1-\alpha}$ marginally results in large confidence sets. In general, the calibration set may not be large enough to accommodate extreme $1-\alpha_{j}$ levels, forced to be closer to 1 as $d$ increases.
@@
We instead model the targets jointly by viewing the target-wise scores as random vectors $S$ that take values in $\mathbb{R}^{d}$. The key question then is how to define the $1-\alpha$ quantile of observed $S$. While ranks and quantiles in $\mathbb{R}$ form the basis of nonparametric statistics (Lehmann et al.,, 1975), a canonical ordering does not exist in $\mathbb{R}^{d}$ for $d>1$. Several notions of multivariate quantiles have been proposed, including ones based on statistical depth (Liu,, 1990; Zuo,, 2003) and geometry (Chaudhuri,, 1996; Hallin et al.,, 2010). We build on the work of Messoudi et al., (2021) that proposed using the copula, a statistical tool for modeling the dependence structure of multivariate random variables Nelsen, (2006). Copulas can be used to estimate the joint cumulative distribution function (CDF) of the scores, $F(s)=\mathbb{P}[S_{1}\leq s_{1},\dotsc,S_{d}\leq s_{d}]$, where $s\in\mathbb{R}^{d}$. The quantile is its generalized inverse, or the $p$-level curve $F^{-1}(p)=\{s\in\mathbb{R}^{d}:F(s)=p\}$. Figure 2 depicts the level curves of the copula (colored contours) fit on correlated scores (gray dots). For $0.91-\alpha=$ 0.9, the $0.9$-level line contains quantiles with smaller norms compared to the independent target approach (green dot), indicating that joint modeling can yield more efficient prediction sets.
@@
As an update to prior work, we estimate $F$ using vine copulas (Bedford and Cooke,, 2002), specifically the nonparametric variant (Nagler et al.,, 2017) for additional expressibility. Vine copulas can scale well to many targets, as they factorize high-dimensional densities into a product of bivariate densities. Applications of copulas for CP have previously been limited to the empirical copula or a small set of parametric families, which can be biased when observations are sparse or the model misspecified (Messoudi et al.,, 2021; Sun and Yu,, 2024). We also demonstrate that a particular copula model allows working with missing-at-random (MAR) target labels in the calibration set.
Having inferred $F$, rather than simply using the plug-in estimate of the quantile, we tailor our analysis to this specific functional by computing the one-step estimator (Le Cam,, 1956) based on its influence function (Hampel,, 1974), best understood as a generalized gradient of the functional with respect to small perturbations of $F$. This debiased estimator specifically targets the quantile estimand of interest and carries better theoretical properties like $\sqrt{n}$-consistency and asymptotic normality (Hines et al.,, 2022).
Contributions: (1) We propose the use of nonparametric vine copulas for flexibly estimating the joint distribution of the non-conformity scores in multivariate conformal calibration (Section 5.1). (2) We propose the semiparametric one-step estimator for debiasing the $1{-}\alpha$ score quantile used to construct the prediction sets (Section 5.2). (3) We prove that our method produces prediction sets with asymptotically exact coverage and present a variant with finite-sample distribution-free validity. (4) We empirically demonstrate the validity and efficiency of our method on various real-world regression problems (Section 6.2).
@@ 2 RELATED WORK
A large body of work aims at extending conformalized quantile regression (Romano et al.,, 2019) to multiple targets (Kong and Mizera,, 2012; Paindaveine and Šiman,, 2011; Feldman et al.,, 2023). There are parallel efforts to make full CP more computationally efficient for some classes of multi-target predictors (Johnstone and Ndiaye,, 2022; Nouretdinov et al.,, 2001). Our focus is to improve the calibration step without additional assumptions on the underlying algorithm.
When there are multiple response variables, a common approach is to design a single scalar score, such as the $L_{2}$ distance between the vector-valued prediction and the label (Kuleshov et al.,, 2018; Johnstone and Cox,, 2021; Messoudi et al.,, 2022; Feldman et al.,, 2023). This implies choosing a target aggregation scheme, however, and the resulting score is sensitive to the relative scales in the predictions and labels across targets. Moreover, the labels for all $d$ targets need to be observed for score evaluation, which may be impractical for large $d$.
Another approach is to conformalize the predictions for each individual target (Neeven and Smirnov,, 2018). Non-conformity scores are evaluated target-wise and the score at the $1-\alpha_{j}$ quantile extracted for each target $j$ such that $\prod_{j}(1-\alpha_{j})=1-\alpha$. If all targets are to be treated equally, we set $1-\alpha_{j}=(1-\alpha)^{1/d}>1-\alpha$ for all $j$. Note that as $d$ increases, $(1-\alpha)^{1/d}$ approaches $1$, leading to inefficient prediction sets. Such extreme levels also require the calibration dataset to be sufficiently large.
If the scores are highly correlated, we can exploit the dependence structure to improve efficiency. Messoudi et al., (2021) pioneered the use of copulas for jointly modeling the distribution of vector scores. Sun and Yu, (2024) applied this idea to multi-step time series forecasting. The empirical copulas used in previous work can be unstable with sparse observations, especially in the upper tails near the quantile level of interest. Messoudi et al., (2021) also uses simple parametric copulas like Gumbel and Gaussian, which offer a smoother fit but can suffer from model misspecification. Copulas are reviewed in Section 4.
@@ 3 Problem setup
A set $\Gamma_{1-\alpha}$ is marginally valid [^1] This is a weaker condition than conditional validity, $\mathbb{P}[Y^{*}\in{\Gamma}_{1-\alpha}(X^{*})|X^{*}]\geq 1-\alpha$. at level $1-\alpha$ if it contains the response $Y^{*}$ with probability at least $1-\alpha$:
@@ 4 BACKGROUND
constructs prediction sets without any assumptions on the data distribution $P_{XY}$ (Vovk et al.,, 2005). The set is guaranteed to be marginally valid for finite samples as long as the data are exchangeable. CP was originally formulated in the transductive setting (Gammerman et al.,, 1998; Saunders et al.,, 1999; Vovk et al.,, 2005). Here we focus on split (inductive) CP (Papadopoulos et al.,, 2002; Lei et al.,, 2018), where the labeled dataset $\{(X^{(i)},Y^{(i)})\}_{i\in\mathcal{I}}$ with index set $\mathcal{I}$ is split into the proper training data $\mathcal{I}_{\rm train}$, used to fit the underlying algorithm ${\hat{f}}$, and the calibration data $\mathcal{I}_{\rm cal}$, used to conformalize it. [^2] We present our framework in terms of split CP for clarity, but it applies to full (transductive) (Vovk et al.,, 2005) and CV (Vovk,, 2015), jackknife (Barber et al.,, 2021) variants as well.
@@
where ${\hat{F}}(s)=\frac{1}{n}\sum_{i\in\mathcal{I}_{\rm cal}}\mathds{1}[S^{(i)}\leq s]$ is the empirical CDF. Applying Equation 2 yields the confidence interval $[{\hat{f}}(X^{*})-Q_{1-\alpha},{\hat{f}}(X^{*})+Q_{1-\alpha}]$. We refer readers to (Shafer and Vovk,, 2008; Angelopoulos and Bates,, 2021) for more details.
@@
Given a random vector $S=[S_{1},\dotsc,S_{d}]$, denote the marginal CDF for each component variable $S_{j}$, $j\in[d]$, as $F_{j}(s_{j})=\mathbb{P}[S_{j}\leq s_{j}]$. The copula $C:[0,1]^{d}\rightarrow[0,1]$ of $S$ is the joint CDF of $U=[F_{1}(S_{1}),\dotsc,F_{d}(S_{d})]$:
@@
Theorem 1 (Sklar’s theorem (Sklar,, 1959)).
@@
The vine copula, or the pair copula construction, can accommodate dependencies of arbitrary complexity when a single standard copula model is inadequate. It is a hierarchical model composed of bivariate copula blocks organized in a sequence of nested trees, also called regular (R-) vines or simply vines Joe, (1994). The main idea is to factorize a $d$-dimensional copula density into a product of ${d(d-1)}/{2}$ bivariate conditional copula densities. Fitting a vine then boils down to choosing the bivariate copula models and the vine structure. The bivariate copulas can be estimated in a nonparametric manner (Chen and Huang,, 2007; Geenens et al.,, 2017; Nagler et al.,, 2017). See Section 9 for more details on nonparametric vine copula estimation.
@@
refers to the inference of nonparametric estimands, or functionals of the true data distribution $P$. Also called targeted or debiased learning, this approach has emerged as a major paradigm in recent years (Van der Laan et al.,, 2011). If the target functional of interest is “smooth” (in ways we will make concrete below), it is sometimes possible to endow estimators with parametric properties like $\sqrt{n}$-consistency and asymptotic normality, even if our estimate of $P$ converges at a slower rate (Murphy and Van der Vaart,, 2000; Van der Laan et al.,, 2011; Kennedy,, 2022; Hines et al.,, 2022).
@@
The second equality follows from restricting ourselves to mean-zero functions for $\psi_{P}$, as the middle expression is insensitive to constant shifts in $\psi_{P}$. When Equation 10 exists and is finite for any “regular” parametric submodel (for which the score $\tilde{P}(Z)/P(Z)-1$ has finite variance), $\psi_{P}$ is called the efficient influence function (EIF) and $\Psi$ is said to be pathwise differentiable (Van Der Vaart,, 1991). The EIF serves as a first-order distributional derivative that characterizes how sensitive $\Psi(P)$ is to changes in $P$ (Robins et al.,, 2017; Kennedy,, 2022).
@@
The one-step estimator (Le Cam,, 1956; Pfanzagl and Pfanzagl,, 1982; Newey et al.,, 1998) is
@@
where $r_{2}$ is the remainder term (Yiu et al.,, 2023). Applying the Riesz representation theorem as in Equation 10 gives
@@
Denoting the estimation error of $\hat{P}$ as $\varepsilon$, we might expect $r_{2}\sim\mathcal{O}(\varepsilon^{2})$. This suggests $\Psi({\hat{P}})+\mathbb{E}_{P}[\psi_{\hat{P}}]$ would be better than the plug-in estimator $\Psi({\hat{P}})$. But $P$ is unknown, of course, so we take the expectation over $\mathbb{P}_{n}$. Figure 3 helps interpret one-step correction as a linear projection Fisher and Kennedy, (2021). See Section 11 for an introduction to one-step estimation.